# Quantifying the Spread of Online Incivility in Brazilian Politics

### v2

@@ 5 Conclusion and Discussion
This finding contrasts with previous studies, which suggest that the right-wing exhibits more uncivil behavior during political campaigns e.g., [46]. A manual post-check of the uncivil posts made by left-aligned and right-aligned influencers revealed that, except for IMP, a large number of uncivil posts are indirect mentions of incivility, meaning that users are discussing or even opposing incivility expressed by others. This phenomenon is more prominent among left-aligned influencers. For instance, ”they are shooting at Indigenous people in the streets (PHAVPR)”, ”I express my full solidarity with an official who was targeted by a misogynistic and cowardly political supporter (HSST)”, ”We are vigilant and monitoring all necessary measures to curb coup-related and anti-democratic acts (THREAT)”  [^2] The examples provided are translated into English and paraphrased to protect user privacy.. This could explain why PHAVPR, HSST, and THREAT are more prevalent during offline violent events led by right-wingers and are primarily associated with left-aligned political influencers. Moreover, determining whether violent metaphors imply actual harm is also challenging. Our results have to be interpreted in light of these limitations, and we advise future studies to exercise caution when applying the classifiers to downstream tasks involving such reverse discourse patterns and metaphors.
@@
The indirect mentions of incivility, especially for dimensions PHAVPR, HSST, and THREAT, can also be found in the discourse of right-aligned political influencers. Even though this may not be called uncivil, indirect mentions of incivility still risk escalating into ”violence for violence”, where users respond to the uncivil behavior of their outgroup in similarly uncivil ways. For instance, we find some direct uncivil expressions from the left-aligned users, such as ”I want to gather everyone from the inauguration and surround Brasília to see who’s stronger (PHAVPR)”, ”An end to these fascist Nazis disguised as Novo (HSST)”, ”It’s time to confront the coup without mediation (THREAT)”. Such cycles can be more detrimental to democratic processes than incivility itself. However, the indirect mentions of incivility can also be beneficial as they draw public attention to stop political violence and enable disadvantaged groups to be seen. Future research could further enhance automated techniques to distinguish between direct and indirect mentions of incivility and assess their respective impacts.
@@
The recent global rise of right-wing populism has often been associated with increased political incivility and violence. In Brazil, this trend was reflected during the Bolsonaro administration (2018–2022). However, growing dissatisfaction with Bolsonaro’s pandemic response and perceived authoritarianism contributed to a left-leaning resurgence, culminating in Lula’s 2022 presidential victory. Lula’s campaign, driven by digital mobilization, has marked a renewed ”pink tide” reminiscent of the early 2000s [31]. This shift may account for the prevalence of both direct and indirect incivility among left-leaning users.
@@
Like many previous studies, this study is not without limitations. For instance, the automatic classification models, while useful, are less accurate than human labeling—a limitation shared by other widely used uncivil detection models like Perspective API [25]. Additionally, we only collect data from political influencers associated with 204 survey users, a rather small sample. However, the sample’s attribute distributions do not differ significantly from the overall survey population, which is representatively sampled from Brazil’s population. Our analysis, therefore, reports how a representative sample experiences uncivil content online. Furthermore, this study focuses on one platform, while right-wing users may be more active on other platforms such as Gab and Telegram. Despite these limitations, this study provides valuable insights into the dynamics of incivility in Brazilian politics and suggests the potential issues of using automatic classifiers detecting incivility.
@@
To protect individual confidentiality, all datasets are securely stored and cannot be shared. The primary Python and R scripts are publicly available via the Harvard Dataverse: https://doi.org/10.7910/DVN/M552GM. Classifiers used to predict multidimensional incivility are available in the GitHub repository: https://github.com/yuanzhang1227/Multidimensional˙Political˙Incivility˙Detection.
6 Acknowledgments
This work has been accepted at ICWSM 2025 and supported in part by funds from the the Swiss National Science Foundation grant 100017_204483.
@@ Appendix A Appendix
A.1 Survey Sample validation
@@
A.2 Identifcation of Political Influencers
@@
    The text include any words related to ideological position leaning (e.g., left, right, liberal, conservative etc.), campaign or social movements slogans (e.g., ”VoltaLula,” ”ForaBolsonaro,” ”Eleicoes2022”), or mentions political/social issues (e.g., ”social welfare”, ”environmental policy”, ”abortion rights”, ”minority rights”).
@@
A.3 Codebook for Multidimentional Incivility Annotation
@@
  • Name-calling (e.g., ”weirdo,” ”traitor,” ”crackpot,” ”thieves”)
  • Aspersions (e.g., ”reckless,” ”stupid,” ”irrational,” ”un-American”)
  • Synonyms for lying (e.g., ”hoax,” ”farce”)
  • Hyperbole (e.g., ”outrageous,” ”heinous”)
  • Words that indicate non-cooperation (e.g., ”polarized,” ”filibuster,” ”inflexible”)
  • Pejorative speak (e.g., ”bellyache,” ”doublespeak”, ”gibberish”)
  • Vulgarity (e.g., ”damn,” ”shit,” ”hell,” ”assholes”)
@@
  • ”The only way to get things done is to take it to the streets and make them listen, by force if necessary.”
@@
  • ”Liberals are less patriotic.”
  • ”Immigrants rely on social benefits.”
  • ”Women are poor drivers.”
@@
  • ”We don’t need elections, we need a strong leader who knows what’s best.”
  • ”The entire electoral process is a sham. Our votes don’t matter.”

### v3

@@ 10 Fl** PO Th** IN Fo** MO Is** IN
This finding contrasts with previous studies, which suggest that the right-wing exhibits more uncivil behavior during political campaigns e.g., (Rega, Marchetti, and Stanziano 2023). A manual post-check of the uncivil posts made by left-aligned and right-aligned influencers revealed that, except for IMP, a large number of uncivil posts are indirect mentions of incivility, meaning that users are discussing or even opposing incivility expressed by others. This phenomenon is more prominent among left-aligned influencers. For instance, "they are shooting at Indigenous people in the streets (PHAVPR)", "I express my full solidarity with an official who was targeted by a misogynistic and cowardly political supporter (HSST)", "We are vigilant and monitoring all necessary measures to curb coup-related and anti-democratic acts (THREAT)"  [^2] The examples provided are translated into English and paraphrased to protect user privacy.. This could explain why PHAVPR, HSST, and THREAT are more prevalent during offline violent events led by right-wingers and are primarily associated with left-aligned political influencers. Moreover, determining whether violent metaphors imply actual harm is also challenging. Our results have to be interpreted in light of these limitations, and we advise future studies to exercise caution when applying the classifiers to downstream tasks involving such reverse discourse patterns and metaphors.
@@
The indirect mentions of incivility, especially for dimensions PHAVPR, HSST, and THREAT, can also be found in the discourse of right-aligned political influencers. Even though this may not be called uncivil, indirect mentions of incivility still risk escalating into "violence for violence", where users respond to the uncivil behavior of their outgroup in similarly uncivil ways. For instance, we find some direct uncivil expressions from the left-aligned users, such as "I want to gather everyone from the inauguration and surround Brasília to see who’s stronger (PHAVPR)", "An end to these fascist Nazis disguised as Novo (HSST)", "It’s time to confront the coup without mediation (THREAT)". Such cycles can be more detrimental to democratic processes than incivility itself. However, the indirect mentions of incivility can also be beneficial as they draw public attention to stop political violence and enable disadvantaged groups to be seen. Future research could further enhance automated techniques to distinguish between direct and indirect mentions of incivility and assess their respective impacts.
@@
The recent global rise of right-wing populism has often been associated with increased political incivility and violence. In Brazil, this trend was reflected during the Bolsonaro administration (2018–2022). However, growing dissatisfaction with Bolsonaro’s pandemic response and perceived authoritarianism contributed to a left-leaning resurgence, culminating in Lula’s 2022 presidential victory. Lula’s campaign, driven by digital mobilization, has marked a renewed "pink tide" reminiscent of the early 2000s (Lampter 2023). This shift may account for the prevalence of both direct and indirect incivility among left-leaning users.
@@
Like many previous studies, this study is not without limitations. For instance, the automatic classification models, while useful, are less accurate than human labeling—a limitation shared by other widely used uncivil detection models like Perspective API (Hosseini et al. 2017). Additionally, we only collect data from political influencers associated with 204 survey users, a rather small sample. However, the sample’s attribute distributions do not differ significantly from the overall survey population, which is representatively sampled from Brazil’s population. Our analysis, therefore, reports how a representative sample experiences uncivil content online. Furthermore, this study focuses on one platform, while right-wing users may be more active on other platforms such as Gab and Telegram. Despite these limitations, this study provides valuable insights into the dynamics of incivility in Brazilian politics and suggests the potential issues of using automatic classifiers detecting incivility.
@@
To protect individual confidentiality, all datasets are securely stored and cannot be shared. The primary Python and R scripts are publicly available via the Harvard Dataverse: https://doi.org/10.7910/DVN/M552GM. Classifiers used to predict multidimensional incivility are available in the GitHub repository: https://github.com/yuanzhang1227/Multidimensional_Political_Incivility_Detection.
Acknowledgments
The authors would like to thank the editor and reviewers for their feedback. This work is supported in part by funds from the the Swiss National Science Foundation grant 100017_204483.
@@ Appendix A Appendix
Survey Sample validation
@@
Identifcation of Political Influencers
@@
    The text include any words related to ideological position leaning (e.g., left, right, liberal, conservative etc.), campaign or social movements slogans (e.g., "VoltaLula," "ForaBolsonaro," "Eleicoes2022"), or mentions political/social issues (e.g., "social welfare", "environmental policy", "abortion rights", "minority rights").
@@
Codebook for Multidimentional Incivility Annotation
@@
  • Name-calling (e.g., "weirdo," "traitor," "crackpot," "thieves")
  • Aspersions (e.g., "reckless," "stupid," "irrational," "un-American")
  • Synonyms for lying (e.g., "hoax," "farce")
  • Hyperbole (e.g., "outrageous," "heinous")
  • Words that indicate non-cooperation (e.g., "polarized," "filibuster," "inflexible")
  • Pejorative speak (e.g., "bellyache," "doublespeak", "gibberish")
  • Vulgarity (e.g., "damn," "shit," "hell," "assholes")
@@
  • "The only way to get things done is to take it to the streets and make them listen, by force if necessary."
@@
  • "Liberals are less patriotic."
  • "Immigrants rely on social benefits."
  • "Women are poor drivers."
@@
  • "We don’t need elections, we need a strong leader who knows what’s best."
  • "The entire electoral process is a sham. Our votes don’t matter."